OpenStudy (anonymous):

the line contains the point (2,-3) and has a slope of (-1/2). Can anyone please help me with this? I need to write an equation in slope-intercept form.

OpenStudy (anonymous):

y-y1=m(x-x1) where (x1,y1) is the any point on the line and m is its slope... so y+3=(-1/2)(x-2) 2y+6=-x+2 2y=-x-4 y=(-1/2)x-2 slope=-1/2 y intercept = -2

OpenStudy (anonymous):

Thanks, I made my mistake on the y+3=(-1/2)(x-2), I wrote y-3. thanks again

OpenStudy (anonymous):

no problem :)